consider th following reaction, initially at equilibrium: N2(g) + 3H2(g) 2NH3(g) + 92 Kj J/mol would the reactants or products have the greatest amount of entropy in this rwacion


1
Expert's answer
2022-06-01T20:50:03-0400

Formation of ammonia is an exothermic reaction

∆S equal for

N2 – 191,5 J/K

H2 – 130,5 J/K

NH3 – 192,6 J/K

N2 + 3H2 = 2NH3

Entropy of reactant =191,5 + 3 *130,5 =583J

Entropy of products = 2 * 192,6 = 385,2 J

ANSWER: the entropy of the reactants is greatest